36) The parameters for the unbalanced Wheatstone Bridge is shown below. The current value of the 1K potentiometer is at 50%. (Center Position) What direction should the potentiometer be turned to achieve a balanced Wheatstone Bridge Vsl 12 V R1 120 R2=100 R3- 220 R3 RL R2 R4 50% A)Less than 50% C) It should remain at 50% B) Greater than 50% D) It cannot be balanced with this Potentiometer 37) The parameters for the unbalanced Wheatstone Bridge is shown below. The current value of the 1K potentiometer is at 50%. (Center Position) What direction should the potentiometer be turned to achieve a balanced Wheatstone Bridge. Vs1 12 V R1 120 R2 220 R3# 330 R1 R3 RL SR2 R4 A) Greater than 50% C) It should remain at 50% B) Less than 50% D) It cannot be balanced with this Potentiometer
Answer #1

Hello,
          Please find the answer attached as under. Please give a thumbs up rating if you find the answer useful! Have a rocking day ahead!

36: The condition for the bridge balance is: (R1/R2) = (R3/R4). Thus, (120/100) = (220/R4), which gives R4 = 183.33 ohms. The 1K pot is at 50% level i.e. 500 ohms. Thus you need to decrease the resistance. For this, you turn the pot less-than-50%.

37: The condition for the bridge balance is: (R1/R2) = (R3/R4). Thus, (120/220) = (330/R4), which gives R4 = 605 ohms. The 1K pot is at 50% level i.e. 500 ohms. Thus you need to increase the resistance. For this, you turn the pot more-than-50%.

NOTE: The voltage supply does not affect the bridge balance condition.

*****************************************************

PS: Please do not forget the thumbs up!​